OpenStudy (anonymous):

In a theather, the ratio of boys to girls is 3: 4 (this means 3 boys for every 4 girls). If there are 24 boys in the theater, how many girls are in the theater ?

OpenStudy (amistre64):

3:4 *8 24:g ; g has to be 32

OpenStudy (anonymous):

How did you get 8 ?

OpenStudy (amistre64):

becasue it says its a ratio of 3:4 and that 3 turns into 24 tells me they scaled it by 8

OpenStudy (amistre64):



OpenStudy (anonymous):

thank-you !

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

we can equate ratio's 3 / 4 =24 / x ; let consider x = number of girls so x=( 24 * 4 / 3) = 32